OpenStudy (anonymous):

i am confused on polynomials: (x^2y+−2x^2y) =-x^2y (the correct answer) x^2-2x^2 = -x, y+y = 2y....am i suppose to add the y's?

OpenStudy (anonymous):

yes they have the same variables (like terms)

OpenStudy (anonymous):

the answer only has one y

OpenStudy (anonymous):

so apparently, you don't do anything to the y's. can someone explain why?

Vocaloid (vocaloid):

the y's are part of the exponent, so you leave them alone

Vocaloid (vocaloid):

think of it this way: 1 - 2 = -1 (x^2y) - 2(x^2y) = -(x^2y)